# Highly Undersampled MRI Reconstruction via a Single Posterior Sampling of Diffusion Models

### v1

@@ Abstract
Incoherent k-space under-sampling and deep learning-based reconstruction methods have shown great success in accelerating MRI. However, the performance of most previous methods will degrade dramatically under high acceleration factors, e.g., 8 or higher. Recently, denoising diffusion models (DM) have demonstrated promising results in solving this issue; however, one major drawback of the DM methods is the long inference time due to a dramatic number of iterative reverse posterior sampling steps. In this work, a Single Step Diffusion Model-based reconstruction framework, namely SSDM-MRI, is proposed for restoring MRI images from highly undersampled k-space. The proposed method achieves one-step reconstruction by first training a conditional DM and then iteratively distilling this model. Comprehensive experiments were conducted on both publicly available fastMRI images and an in-house multi-echo GRE (QSM) subject. Overall, the results showed that SSDM-MRI outperformed other methods in terms of numerical metrics (PSNR and SSIM), qualitative error maps, image fine details, and latent susceptibility information hidden in MRI phase images. In addition, the reconstruction time for a 320320 brain slice of SSDM-MRI is only 0.45 second, which is only comparable to that of a simple U-net, making it a highly effective solution for MRI reconstruction tasks.
MRI Acceleration, Single-Step Diffusion Model (SSDM), QSM Acceleration, Image Reconstruction
@@ 1 Introduction
Incoherent undersampling combined with compressed sensing (CS) reconstruction algorithms has been substantially studied to accelerate MRI scans, by solving an ill-posed inverse problem. Many traditional iterative reconstruction algorithms have been developed since the pioneering work of Lustig [3, 4], such as image sparsity-based CS methods [3, 4, 5, 6, 7] or low-rankness models [8, 9, 10]. However, the performance of these algorithms will substantially decrease at high acceleration factors, leading to image blurring or residual artifacts. In addition, they are usually computationally expensive and require manual hyperparameter tuning [11].
@@
Deep learning (DL) algorithms have been increasingly popular as an alternative to conventional iterative methods for MRI acceleration. Many end-to-end supervised deep neural networks [12, 13, 14, 15], such as U-net [16] based methods, have been proposed for MRI reconstruction recently, demonstrating better artifact removal and improved fine details, compared with traditional iterative methods. However, the generalization capability of these methods is limited, and their performance will degrade when applied to data out of the training distribution. Whereas, unrolling-based or model-based deep neural network [17, 18] is another framework incorporating physical models in deep neural network designs, demonstrating better results as well as generalization capability. However, these two frameworks will both decrease their performance when applied to highly undersampled k-space data (e.g., 8 or higher).
Recently, diffusion models (DM)[19, 20, 21, 22, 23] or noise-conditioned score networks are increasingly popular for MRI reconstruction at high acceleration factors [24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. This new framework holds great potential for achieving ultra-high k-space acceleration factors, as long as a desirable generative model is successfully trained, which allows high-quality MRI synthesis from only random noise. However, a dramatic number of iterative reverse posterior sampling steps are usually necessary for DM to generate reliable results, which will accumulate reconstruction times, making DM-based MRI reconstruction framework much slower than previous DL methods, e.g., end-to-end network-based methods [12, 13, 14].
@@
In this work, a Single Step Diffusion Model-based method, namely SSDM-MRI, is designed to reconstruct MRI images from highly subsampled k-space in just one posterior sampling step, thus enabling much faster computation than conventional DM-based methods. The proposed method was developed by progressively distilling a pre-trained conditional DM, which was obtained through a supervised learning scheme to enhance the stability and controllability. Comprehensive experiments were conducted on both the public fastMRI brain and knee datasets [34] and an in-house complex-valued brain dataset to compare the proposed SSDM-MRI with multiple established MRI reconstruction methods, including both iterative algorithms and deep learning frameworks. The performances of SSDM-MRI at different acceleration rates were also investigated. In addition, we also evaluated SSDM-MRI’s performance on Quantitative Susceptibility Mapping (QSM) acceleration task using the in-house brain dataset.
@@ 2.1 Problem Formulation
where represents the undersampled k-space measurements and is the underlying image, and ; is the measurement noise, and represents the system forward model, with being the undersampling mask and as the Fourier transform. Equation (1) is also known as the general forward model of MRI reconstruction, serving as a training data generator in many supervised learning frameworks [35, 14, 36].
@@ 2.2 Conventional Iterative and Deep Learning Algorithms
where ; is the magnitude images, and is the phase signals; is a loss function enforcing data fidelity, e.g., the commonly used l2 norm; is a manually-designed regularization term, which can be a sparsifying transformation (e.g., Total Variation and Wavelet Transform), and is a weighting hyperparameter. In most previous studies [3, 37, 38, 39], the phase signal is simply estimated from the low-frequency part of k-space to simplify this problem; however, a complex phase regularizer is also necessary for phase-based advanced applications [40, 41, 15], e.g., Quantitative Susceptibility Mapping (QSM).
Deep neural networks, due to their excellent performances in complex feature representation capability compared with manually-designed regularizers (i.e., in Eq. (2)), have demonstrated improved reconstruction performances. End-to-end neural networks [12, 13, 14] and deep unrolling networks [17, 18] are two very commonly used frameworks in CS-MRI. The former can be briefly described as follows:
where is the learnable parameters in the deep network , and is the loss function adopted for network training; represents the techniques used to address the potential overfitting problems in deep and large models. In this scheme, end-to-end networks tried to directly learn the non-linear mapping between training inputs, which is usually the zero-filling reconstructions (i.e., ), and training labels (). Whereas deep unrolling networks, on the other hand, are usually constructed by iteratively stacking physical-model modules and neural networks. The network part usually tries to learn the mapping between latent variables in conventional iterative algorithms (i.e., learning the proximal gradient descent process in ISTA-Net [18]).
@@ 2.3 Diffusion Models for MRI Reconstruction
Diffusion models (DM)[19, 21, 22] or the noise-conditioned score networks (NCSN)[20, 23] are new types of generative models that are capable of synthesizing high-quality images via gradually denoising Gaussian noise. It involves a forward noise adding process to convert the original image into pure Gaussian noise, and an iterative backward sampling process to recover the original images in consecutive steps. DM is naturally suitable for solving inverse problems and has shown great potential in MRI reconstructions [24, 25, 26, 27, 28, 29, 30, 31].
Suppose that the desired MRI images follow a specific unknown distribution (i.e., ), then the reconstruction of MRI images () from subsampeld k-space () can be formulated as a posterior sampling process , and the key to obtain reliable samples from this conditional distribution in the DM backward process is to estimate the corresponding score function [20, 42] (i.e., the noise estimation in DM). One can train a deep neural network to estimate this conditional score () in a supervised or unsupervised manner.
@@ 2.4 Progressive Diffusion Model Distillation
One major drawback of DM-based methods is the slow sampling (backward) process, which usually takes a large number of posterior sampling steps. Distillation strategies were commonly employed to accelerate the posterior sampling process in many previous works [23, 43], among which progressive distillation [44] has been one of the most effective techniques. It iteratively reduces the number of necessary sampling steps by distilling a teacher diffusion model into a faster student model, which then becomes a new teacher for the next iterative distillation. Every distillation reduces the number of necessary sampling steps from steps to steps.
@@ 3.1 Overall Framework of SSDM-MRI
The proposed SSDM-MRI takes the zero-filling reconstructions as well as the Gaussian noise as inputs (concatenated along the channel dimension), and generates desired MRI images as outputs, as demonstrated in Fig. 1. To achieve robust image reconstruction with only one posterior sampling, a conditional denoising diffusion probability model (DDPM) [45] was first trained using supervised learning, and it was then progressively distilled to speed up the sampling process, decreasing necessary reverse sampling steps to only one step. Data consistency [46, 15] was also performed to further refine the DM-synthesised images to impose the k-space data fidelity, as described in Eq. (4) below:
@@
3.2 Diffusion Model Pretraining and Distillation
@@
The pre-trained conditional DM was obtained by training a self-tailored residual U-net () on a large number of paired datasets () simulated based on Eq. (1), as detailed in Algorithm 1. Overall, the U-net takes both and zero-filling reconstructions () as its major inputs (concatenated along the channel dimension). Similar to most previous U-net models, this adopted U-net consists of an encoding and a decoding path, each consisting of three modules, and each module contains three convolutional blocks with residual connections from the block input to the output. Three pooling layers and three corresponding unpooling layers were included for multiscale or multi-resolution learning. Self-attention operations were only carried out on the latent features of the lowest resolution due to memory constraints. Overall, the network consists of 52 convolutional layers (kernel size: 33), 3 max-pooling layers (kernel size: 22), 3 transposed convolutional layers (kernel size: 22), 51 batch-normalization layers, 77 rectified linear units (ReLUs), 3 concatenations, 25 residual connections, and 6 self-attention operations. The network also takes the noise scheduler () as its third input, which was encoded using a simple thee-layer multi-layer perceptron, inspired by previous works [45] focusing on conditional generation.
@@
In this work, the network is designed to predict a latent variable , which is better suited for distillation [44], considering the fact that the standard noise prediction parameterization used in DDPM is not well suited for distillation, particularly as the signal-to-noise ratio (SNR) approaches zero, where small errors in noise prediction can be significantly amplified, leading to unstable training and degraded performance. Moreover, when distilling down to very few sampling steps, the connection between noise prediction and data reconstruction becomes unreliable. Using the v-diffusion parameterization, the denoising process turns into , which better preserves stable supervision signals during distillation and enables more efficient sampling. Following this modification, progressive distillation is then carried out iteratively to achieve robust image reconstruction to achieve one-step reconstruction. In each distillation iteration, the student model is optimized to halve the overall sampling steps to by learning to replicate the behavior of two sampling steps of the teacher model in just one step, as detailed in Algorithm 2.
@@ 3.3 SSDM-MRI for multi-coil data and QSM acceleration
Similar to our previous work [15], the reconstruction of SSDM-MRI for multi-coil data can be carried out as follows: (1) zero-filling reconstruction was first performed for each receiver channel, (2) then, nm the coil sensitivity maps were estimated using the POEM method for coil combination [47], (3) next, the combined images were fed into SSDM-MRI network (without data consistency) for initial reconstruction, (4) then the initial reconstructions were multiplied by the estimated coil sensitivities to convert back into a multi-channel k-space, (5) finally, data consistency was conducted for every coil to obtain images for each coil channel. Finally, these multi-coil images were re-combined using the POEM method [47].
Although the proposed SSDM-MRI is designed for 2D MRI reconstructions, it can be easily transferred to advanced 3D MRI-based applications in a slice-by-slice manner with proper 1D Fourier Transform preprocessing and k-space data consistency, if the acquisition follows the strategy proposed in our previous QSM acceleration work [15]. iQSM+ [48] was adopted for QSM reconstruction from reconstructed MRI phase images.

### v2

@@ Abstract
Incoherent k-space undersampling and deep learning-based reconstruction methods have shown great success in accelerating MRI. However, the performance of most previous methods will degrade dramatically under high acceleration factors, e.g., 8 or higher. Recently, denoising diffusion models (DM) have demonstrated promising results in solving this issue; however, one major drawback of the DM methods is the long inference time due to a dramatic number of iterative reverse posterior sampling steps. In this work, a Single Step Diffusion Model-based reconstruction framework, namely SSDM-MRI, is proposed for restoring MRI images from highly undersampled k-space. The proposed method achieves one-step reconstruction by first training a conditional DM and then iteratively distilling this model four times using an iterative selective distillation algorithm, which works synergistically with a shortcut reverse sampling strategy for model inference. Comprehensive experiments were carried out on both publicly available fastMRI brain and knee images, as well as an in-house multi-echo GRE (QSM) subject. Overall, the results showed that SSDM-MRI outperformed other methods in terms of numerical metrics (e.g., PSNR and SSIM), error maps, image fine details, and latent susceptibility information hidden in MRI phase images. In addition, the reconstruction time for a 320320 brain slice of SSDM-MRI is only 0.45 second, which is only comparable to that of a simple U-net, making it a highly effective solution for MRI reconstruction tasks.
MRI Reconstruction, Single-Step Diffusion Model (SSDM), QSM Acceleration, Image Reconstruction
@@ 1 Introduction
Incoherent undersampling combined with compressed sensing (CS) reconstruction algorithms has been substantially studied to accelerate MRI scans, by solving an ill-posed inverse problem. Many traditional iterative reconstruction algorithms have been developed since the pioneering work of Lustig [3, 4], such as image sparsity-based CS methods [3, 4, 5, 6, 7, 8, 9, 10] or low-rankness models [11, 12, 13, 14]. However, the performance of these algorithms will substantially decrease at high acceleration factors, leading to image blurring or residual artifacts. In addition, they are usually computationally expensive and require manual hyperparameter tuning [15].
@@
Deep learning (DL) algorithms have been increasingly popular as an alternative to conventional iterative methods for MRI acceleration. Many end-to-end supervised deep neural networks [16, 17, 18, 19, 20, 21, 22], such as U-net [23] based methods, have been proposed for MRI reconstruction recently, demonstrating better artifact removal and improved fine details, compared with traditional iterative methods. However, the generalization capability of these methods is limited, and their performance will degrade when applied to data out of the training distribution. Whereas, unrolling-based or model-based deep neural network [24, 25, 26, 27] is another framework incorporating physical models in deep neural network designs, demonstrating better results as well as generalization capability. However, these two frameworks will both decrease their performance when applied to highly undersampled k-space data (e.g., 8 or higher).
Recently, diffusion models (DM)[28, 29, 30, 31, 32] or noise-conditioned score networks are increasingly popular for MRI reconstruction at high acceleration factors [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43]. This new framework holds great potential for achieving ultra-high k-space acceleration factors, as long as a desirable generative model is successfully trained, which allows high-quality MRI synthesis from only random noise. For example, Score-MRI [33] successfully learned the time-dependent score functions for high-quality MRI images generation, achieving promising reconstruction quality by iteratively refining undersampled data through stochastic denoising and data-consistency enforcement (at least hundreds of reverse sampling steps required). DiffINR [34] proposed a novel posterior sampler for diffusion models using Implicit Neural Representation (INR), which demonstrated better data consistency and more accurate and stable reconstructions at high acceleration rates. It also reduced the necessary number of reverse sampling steps from at least 1000 steps using Score-MRI to only 250 steps using DiffINR. A Predictor-Projector-Noiser (PPN) method [35] was also developed based on Denoising Diffusion Implicit Models (DDIM) [32] for controllable and efficient diffusion sampling by repetitively enforcing data fidelity during each generative step. A zero-shot Bi-level Guided Diffusion Model (BGDM) [36] was developed to solve the inverse problems in medical imaging by firstly adopting an inner-level conditional posterior to obtain an initial prediction and then reinforcing the measurement consistency in the outer-level proximal optimization. High-Frequency Space Diffusion Model (HFS-SDE) [37] was another framework that was able to accelerate the convergence of the reverse process by confining the diffusion process to high-frequency k-space regions while keeping low-frequency components deterministic, showing improved structural fidelity. SPIRiT-Diffusion [38] chose to incorporate a SPIRiT-based self-consistency prior into the diffusion process, enabling physics-guided reconstruction that enhances robustness under high acceleration factors. More recently, a Come-Closer-Diffuse-Faster (CCDF) [40] framework was proposed to accelerate the reverse diffusion process of conditional DM models by starting from a forward diffusion point from the initial estimates instead of from pure noise, and it significantly reduces the number of reverse steps to only 20 steps compared to 1000 of Score-MRI. Despite these improvements, a dramatic number of iterative reverse posterior sampling steps are still necessary for these methods to generate reliable and high-quality MRI images, which will accumulate reconstruction times (increasing with more numbers of reverse sampling steps), making the DM-based MRI reconstruction framework much slower than previous DL methods, e.g., end-to-end network-based methods [17, 18, 19].
@@
In this work, a Single Step Diffusion Model-based method, namely SSDM-MRI, is designed to reconstruct MRI images from highly subsampled k-space in just one posterior sampling step, thus enabling much faster computation than conventional DM-based methods. The proposed method was developed by progressively distilling a pre-trained conditional DM four times using our proposed iterative selective distillation algorithm, which was performed collaboratively with a shortcut reverse sampling strategy to enhance the stability and controllability. Comprehensive experiments were conducted on both the public fastMRI brain and knee datasets [44] and an in-house complex-valued brain dataset to compare the proposed SSDM-MRI with multiple established MRI reconstruction methods, including both iterative algorithms and deep learning frameworks. The performances of SSDM-MRI at different acceleration rates were also investigated. In addition, we also evaluated SSDM-MRI’s performance on Quantitative Susceptibility Mapping (QSM) acceleration task using the in-house multi-echo brain dataset, and R2* images were also calculated to verify the performance of the proposed SSDM-MRI in multi-echo magnitude imaging contexts.
@@ 2.1 Problem Formulation
where represents the undersampled k-space measurements and is the underlying image, and ; is the measurement noise, and represents the system forward model, with being the undersampling mask and as the Fourier transform. Equation (1) is known as the general forward model of MRI reconstruction, and it can be used as a data generator to simulate subsampled k-space data from the fully-sampled MRI images, which is popular in many supervised learning frameworks [16, 19, 22]. In this work, we also adopted Equation (1) to generate training datasets for network training.
@@ 2.2 Conventional Iterative and Deep Learning Algorithms
where ; is the magnitude images, and is the phase signals; is a loss function enforcing data fidelity, e.g., the commonly used l2 norm; is a manually-designed regularization term, which can be a sparsifying transformation (e.g., Total Variation and Wavelet Transform), and is a weighting hyperparameter. In most previous studies [3, 5, 6, 21, 27], the phase signal is simply estimated from the low-frequency part of k-space to simplify this problem; however, a complex phase regularizer is also necessary for phase-based advanced applications [10, 20], e.g., Quantitative Susceptibility Mapping (QSM).
Deep neural networks, due to their excellent performances in complex feature representation capability compared with manually-designed regularizers (i.e., in (2)), have demonstrated improved reconstruction performances. End-to-end neural networks [17, 18, 19] and deep unrolling networks [24, 25] are two very commonly used frameworks in CS-MRI. The former can be briefly described as follows:
where and denote the training sample pair of sub-sampled k-space data and the corresponding fully-sampled image in the training dataset, respectively, and is the learnable parameters in the deep network , and is the loss function adopted for network training; represents the techniques used to address the potential overfitting problems in deep and large models. In this scheme, end-to-end networks tried to directly learn the non-linear mapping between training inputs, which is usually the zero-filling reconstructions (i.e., ), and training labels (). Whereas deep unrolling networks, on the other hand, are usually constructed by iteratively stacking physical-model modules and neural networks. The network part usually tries to learn the mapping between latent variables in conventional iterative algorithms (i.e., learning the proximal gradient descent process in ISTA-Net [25]).
@@ 2.3 Diffusion Models for MRI Reconstruction
Diffusion models (DM)[28, 30, 31] or the noise-conditioned score networks [29, 32] are new types of generative models that are capable of synthesizing high-quality images via gradually denoising Gaussian noise. It involves a forward noise adding process to convert the original image into pure Gaussian noise, and an iterative backward sampling process to recover the original images in consecutive steps. DM is naturally suitable for solving inverse problems and has shown great potential in MRI reconstructions [34, 41, 33, 37, 42, 43, 38, 39].
Suppose that the desired MRI images follow a specific unknown distribution (i.e., ), then the reconstruction of MRI images () from subsampled k-space () can be formulated as a posterior sampling process , and the key to obtain reliable samples from this conditional distribution in the DM backward process is to estimate the corresponding score function [29, 45] (i.e., the noise estimation in DM). One can train a deep neural network to estimate this conditional score () in a supervised or unsupervised manner.
@@ 2.4 Progressive Diffusion Model Distillation
One major drawback of DM-based methods is the slow (reverse) sampling process, which usually takes a large number of posterior sampling steps. Distillation strategies were commonly employed to accelerate the posterior sampling process in many previous works [32, 46], among which progressive distillation [47] has been one of the most effective techniques. It iteratively reduces the number of necessary sampling steps by distilling a teacher diffusion model into a faster student model, which then becomes a new teacher for the next iterative distillation. Every distillation reduces the number of necessary sampling steps from steps to steps.
2.5 CCDF: Accelerating the Sampling of Conditional Diffusion Models
Come-Closer-Diffuse-Faster (CCDF) [40] is an efficient and effective framework for accelerating the reverse sampling of pretrained conditional diffusion models for inverse problems. It is shown that there exists a shortcut reverse sampling path for a pretrained conditional DM, so that the inference can start from a closer point and achieve improved final results in much fewer sampling steps. The key design is to start the inference of DMs with a forward diffusion of the initial estimates of the desired images, instead of from pure Gaussian noise. As report in the original work[40], 20 reverse sampling steps are enough for CCDF to obtain more accurate reconstructions than Score-MRI [33] of 1000 reverse steps.
@@ 3.1 Overall Framework of SSDM-MRI
The proposed SSDM-MRI was constructed by two steps: (1), training a conditional denoising diffusion probability model (DDPM) [28] in a supervised manner, which required reverse sampling steps to generate clean images () from pure Gaussian noise () and the zero-filling image (, i.e., the generation condition), as shown in Fig. 1(a), and (2), distilling the pretrained DM using the proposed iterative selective distillation algorithm, which was performed in pairs with the shortcut reverse sampling strategy. The proposed selective distillation algorithm and the shortcut reverse sampling strategy are detailed in Subsection 3.3.
As shown in the bottom panel of Fig. 1(b), the proposed SSDM-MRI takes the zero-filling image () and a latent image ( detailed in Subsection 3.3) forward diffused from the zero-filling image as inputs (concatenated along the channel dimension), and generates desired MRI images as outputs. To achieve more robust and stable image reconstruction, a projection-based data consistency [21, 20] was also performed to further refine the DM-synthesized images to impose the k-space data fidelity, as described in (4) below:
@@
3.2 Conditional Diffusion Model Pretraining
@@
The pre-trained conditional DM was obtained by training a self-tailored residual U-net () on a large number of paired datasets () simulated based on (1), as detailed in Algorithm 1. Overall, the U-net takes both and zero-filling reconstructions () as its major inputs (concatenated along the channel dimension). Similar to most previous U-net models, this adopted U-net consists of an encoding and a decoding path, each consisting of three modules, and each module contains three convolutional blocks with residual connections from the block input to the output. Three pooling layers and three corresponding unpooling layers were included for multiscale or multi-resolution learning. Self-attention operations were only carried out on the latent features of the lowest resolution due to memory constraints. Overall, the network consists of 52 convolutional layers (kernel size: 33), 3 max-pooling layers (kernel size: 22), 3 transposed convolutional layers (kernel size: 22), 51 batch-normalization layers, 77 rectified linear units (ReLUs), 3 concatenations, 25 residual connections, and 6 self-attention operations. The network also takes the noise scheduler () as its third input, which was encoded using a simple thee-layer multi-layer perceptron, inspired by previous works [48] focusing on conditional generation.
@@
In this work, the network is designed to predict a latent variable , which is better suited for distillation [47], considering the fact that the standard noise prediction parameterization used in DDPM is not well suited for distillation, particularly as the signal-to-noise ratio (SNR) approaches zero, where small errors in noise prediction can be significantly amplified, leading to unstable training and degraded performance. Using the v-diffusion parameterization, the denoising process turns into , which better preserves stable supervision signals during distillation and enables more efficient sampling. Following this modification, the proposed iterative selective distillation is then carried out, and in each distillation iteration, the student model is optimized to replicate the behavior of two sampling steps of the teacher model in just one step, as detailed in Algorithm 2 (line 7-21).
3.3 Iterative Selective Distillation and Shortcut Reverse Sampling
Fig. 1(b) illustrates the main concepts of the proposed iterative selective distillation algorithm (the green block) and a paired shortcut reverse sampling strategy (the orange block), which were conducted synergistically to reduce the reverse sampling steps of the pretrained DM from the original steps to only one step (blue block). The implementation details of the proposed iterative selective distillation are depicted in Algorithm 2, which is inspired by and modified from the progressive distillation algorithm [47]. The major improvement is that the current selective distillation strategy only carries out the distillation process on the latter reverse sampling steps (from to clean image , i.e., the green circles in the dashed boxes in Fig. 1(b)), as also indicated in line 3 and 6 of Algorithm 2, while the conventional progressive distillation strategy is performed on the entire reverse sampling path (all reverse sampling steps). Furthermore, the proposed Algorithm 2 also incorporated a DC operation on the latent results of the teacher and student models, i.e., and in line 18, and a second loss was introduced on the DC results in line 20.
As the first half reverse sampling path (from pure noise to ) is ignored during the distillation, the sampling method of the proposed SSDM-MRI was updated correspondingly in this work. Inspired by CCDF [40], a shortcut reverse sampling strategy was adopted to start the reverse sampling of the pretrained DM from a latent point instead of pure noise (), which was calculated by a forward diffusion on the zero-filling reconstruction (the blue dashed line in Fig. 1(b)) using the following equation:
where is the zero-filling reconstructions from the subsample k-space data (); as defined in 5. is a latent point on the reverse sampling path of the pretrained DM, and it is empirically set as in this work.
Combining the iterative selective distillation and shortcut reverse sampling, the numbers of the reverse sampling steps for the distilled models after the distillation stage is reduced from the original steps to steps only, which is much more efficient than the conventional progressive distillation, where the resulting model of distillation will need steps for high quality image generation. Therefore, for the pretrained conditional DM in the previous section (), only four iterative distillations (i.e., set in Algorithm 2) are enough to obtain a single-step model for the proposed SSDM-MRI. In addition, as the student model’s performance would decrease more with an increasing number of distillations, compared with the pretrained model, there would be fewer performance declines using the proposed iterative selective distillation and shortcut reverse sampling framework. Further, the proposed design is intrinsically beneficial to improve the model’s performance because the student model during model distillation now only needs to learn half of the dynamics in the reverse diffusion path, rather than the entire reverse path from the pure Gaussian noise to clean images.
@@ 3.4 SSDM-MRI for Multi-Coil Data and QSM Acceleration
As shown in Fig. 2, the reconstruction of SSDM-MRI for multi-coil data can be carried out using the following pipeline: (1) zero-filling reconstructions was first performed for each receiver channel, (2) then, the coil sensitivity maps were estimated using Walsh’s (SVD-based) method [49] for coil combination, because it can estimate the coil sensitivities in image-domain without explicitly specifying ACS lines, (3) next, the combined images were fed into SSDM-MRI network (without data consistency) for initial reconstruction, (4) then the initial reconstructions were multiplied by the estimated coil sensitivities to convert back into a multi-channel k-space, (5) next, data consistency was conducted for every coil to obtain images for each coil channel, (6) finally, these multi-coil images were re-combined using Walsh’s (SVD-based) method [49].
Although the proposed SSDM-MRI is designed and trained for 2D MRI image reconstructions, it can be easily transferred to advanced 3D MRI-based applications in a slice-by-slice manner, following the strategy proposed in our previous QSM acceleration work [20], where the 3D subsampled k-space volumes are first converted into 2D k-space slices after a 1D Fourier Transform along the fully-sampled (frequency-encoding) direction before being fed into the networks for reconstructions. 3D QSM images can then be calculated from the reconstructed MRI phase images, using best path phase unwrapping [50, 51], RESHARP-based background removal [52], and COSMOS dipole inversion [53].
In this work, the 2D subsampling patterns involved were only applied to 3D k-space volumes to make sure that the readout (frequency-encoding) direction is always fully sampled to meet the MRI acquisition convention.